OpenStudy (anonymous):

Find conditions on a and b such that the following system of linear equations has a. no solution b. a unique solution c. infinitely many solutions

OpenStudy (anonymous):

what r the equations?

OpenStudy (anonymous):

x+ y +3z=2

OpenStudy (anonymous):

sorry i mean

OpenStudy (anonymous):

I don't get the questions.

OpenStudy (anonymous):

x+ y +3z=2 x+2y+4z=3 x+3y+az=b

OpenStudy (anonymous):

thats the equations

OpenStudy (anonymous):

make it as AX=B where A is a 3 X 3 matrix. X= A^-1 B